OpenStudy (anonymous):

Sandy cut out 2 squares and 7 circles from a piece of cardboard. The length of the sides of each square was x inches, and radius of each circle was y inches. The expression below shows how many more square inches of cardboard Sandy used for making the circles than the squares: 7πy2 - 2x2 What does the term 7πy2 in the expression represent? The total area of 7 circles The total area of 22 circles The total area of all the shapes The total area of all the squares

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I think it's A

OpenStudy (phi):

can you explain what you're thinking?

OpenStudy (anonymous):

Well it says 7πy2 and I think the only way it could be 22 circles is if it was 22πy2

OpenStudy (anonymous):

Sorry it took me forever to reply, I'm working on the other problems and I forgot about this one.

OpenStudy (phi):

yes, your thinking is good. You should mention that the area of a circle is \[ \pi r^2\] when the radius is "y" (as in this problem), the area of one circle is \[ \pi y^2 \] and 7 circles will have 7 times this area: \[ 7 \pi y^2 \]